OpenStudy (anonymous):

solve for x. 3/x=x/32

OpenStudy (zzr0ck3r):

multiply both sides by x and tell me what you get

OpenStudy (anonymous):

you mean 3x32?

OpenStudy (anonymous):

or x*x= x squared?

OpenStudy (zzr0ck3r):

we have \[\frac{ 3 }{ x }\] = \[\frac{x}{32}\]

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

so multiply both sides by x and we have 3 = x^2/32 right?

OpenStudy (anonymous):

yes but what about the 3?

OpenStudy (zzr0ck3r):

now muliply both sides by 32 now we have 32*3 = x^2 right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

96

OpenStudy (zzr0ck3r):

ok so now we have x^2 = 96 can you solve for x?

OpenStudy (zzr0ck3r):

note: sqrt(96) = sqrt(16*6) = sqrt(16)*sqrt(6) = ?

OpenStudy (zzr0ck3r):

@meylispena we are not here to give answers, please let them participate. its not right also

OpenStudy (anonymous):

I deleted it, I'm sorry for the inconvenience.

OpenStudy (anonymous):

is the answer \[4\sqrt{6}\] ?

OpenStudy (zzr0ck3r):

what is sqrt(4)?

OpenStudy (anonymous):

oh I didn't even read your replies before I wrote that, my internet shut down for a minute.

OpenStudy (anonymous):

2

OpenStudy (zzr0ck3r):

+-2

OpenStudy (anonymous):

0?

OpenStudy (zzr0ck3r):

so sqrt(16)*sqrt(6) = +- 4*sqrt(6)

OpenStudy (zzr0ck3r):

there are two answers....